# Activated solids: Spontaneous deformations, non-affine fluctuations, softening, and failure

### v1

@@ Abstract
We investigate spontaneous deformations in activated solids by analyzing crystal deformations through non-affine fluctuations. A scaling law for non-affinity is derived, showing it scales quadratically with activity and inversely with density. At high activity, the non-affine parameter diverges, signaling defect proliferation and melting to a hexatic phase. This softening is marked by reduced shear modulus, solid order, and hexatic order. At a higher activity, the hexatic melts to a fluid. We propose locally tuned activity to control non-affinity, demonstrated through numerical simulations. Our predictions are testable in active colloids performing persistent random motion.
@@ I Introduction
The study of active solids is inspired by biological systems like cell assemblies, tissues, and bacterial biofilms, as well as artificial ones such as robotic solids [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. These systems are out of equilibrium at microscopic scales, with elements that locally consume and dissipate energy, generating propulsion or stress and breaking detailed balance. Unlike crystalline solids, which break continuous translational and rotational symmetries in equilibrium, amorphous solids can be stabilized by compression and are governed by local force and torque balance [16, 17, 18, 19]. In active solids maintaining such balance is essential for stability, as any departure can lead to deformation and flow.
Drawing inspiration from nature, various artificial micro- and nano-swimmer systems have been developed over last decades [20, 21]. These self-propelled colloids move by harnessing ambient energy – optical, thermal, electrical, or chemical – through an auto-catalytic process. Their behavior can be modeled by the active Brownian particle (ABP) framework, where particles move at a constant speed in their direction of motion, with random reorientations. This paper examines the impact of such activity on a sterically stabilized solid.
Equilibrium solids display low-energy phonon modes, described by displacement fields and elasticity theory. At higher excitations, topological defects, manifested as singularities in the displacement field, emerge. This framework not only captures the behavior of plastic deformation and yielding in various solids [22, 23, 24], but also offers a robust description of two-dimensional solid melting from quasi-long ranged ordered to disordered phase within the BKTHNY theory, driven by the proliferation of topological defects [25, 26, 27, 28, 29]. While the displacement field approach can be applied to amorphous solids with limited success [30], the absence of broken translational symmetry complicates distinguishing smooth from singular displacements [31, 32]. Instead, atomic configurations with localized, non-singular displacements – both proposed [33, 30] and observed experimentally [34] – are thought to function analogously to lattice defects in these materials.
A generalized approach to classifying atomic displacements in solids have been developed over the last decade [35, 36, 37, 38, 39, 40, 41, 42, 43]. Building on concepts used to identify local plastic events in structural glasses [17], it has been shown that atomic displacements within a coarse-graining region can be projected onto two orthogonal subspaces. The affine component represents homogeneous linear transformations, such as isotropic expansion, shear strains, and rotations, while the nonaffine component captures the residual displacements. These components are linearly independent [36]. Nonaffine modes have been linked to defect precursors in various lattice structures [42], bridging the projection formalism with conventional plasticity theories. This approach has also addressed fundamental questions in solid rigidity [41] and dynamic yield strength in nanoscale crystals [44]. Experimentally, nonaffine rearrangements have been observed to correlate with structural anisotropy in metallic glasses [45] and colloidal crystals [46].
In this paper, we explore spontaneous deformations in active solids using the concepts discussed above. The study of active matter [21, 20] has revealed novel non-equilibrium phenomena, such as flocking [47, 48], pattern formation and instabilities in active polar or apolar nematics [49, 50, 51, 52, 53], and motility-induced phase separation in active Brownian particles [54, 55, 56, 57], with their mechanisms understandable as low density phenomena [58, 59, 55, 60]. Experimental realizations involve, apart from natural flocks and swarms, rely on active colloids, active granular or robotic systems such as hexbugs [20, 61, 62, 63, 64, 65, 66, 67, 12, 13]. At high densities, active systems can exhibit liquid crystalline, jammed, or crystalline states [68, 60, 69, 70, 71, 72, 15]. Recent studies also highlight extreme spontaneous deformations in active crystals [73]. In this study, we characterize crystal deformations using non-affine fluctuations, deriving a scaling law in terms of activity $\Lambda$ and density. Our scaling argument predicts a $\Lambda^{2}$ dependence, confirmed over three decades through detailed numerical simulations – one of the key achievements of this paper. The non-affine parameter diverges at large $\Lambda$, signaling defect proliferation and activated melting of the solid to a hexatic phase, marked by vanishing shear modulus, solid order, and eventually the hexatic order. The softening and melting are further characterized analyzing the proliferation of topological defects. Finally, we propose locally tuned activity as a means to control non-affinity and the mechanical properties of a solid.
@@ II Model
We investigate the behavior of an equilibrium solid at high density ($\rho\sigma^{2}\geq 1$) under self-propulsion, modeled by Active Brownian Particles (ABP). Each particle moves with self-propulsion speed $v_{0}$ in a direction $\hat{\mathbf{n}}_{i}={\cos\theta_{i}(t),\sin\theta_{i}(t)}$, with the angle $\theta_{i}(t)$ evolving via rotational diffusion. The dynamics of the $i$-th particle is governed by the following coupled Langevin equations:
$d{\bf r}_{i}(t)$ $=v_{0}\hat{\bf n}_{i}(t)dt-\mu\nabla_{i}\sum_{j\epsilon R_{i}}U(r_{ij})dt+\sqrt{2D_{t}dt}\,d{\bf B}_{i}(t)$
$d{\theta_{i}}(t)$ $=\sqrt{2D_{r}dt}\,dB^{r}_{i}(t),$ (1)
where $D_{t}$ and $D_{r}$ are the translational and rotational diffusivities, and $d\mathbf{B}_{i}(t)$ and $dB_{i}^{r}(t)$ are independent and univariate Gaussian white noise terms with mean zero. The Euler-Maruyama method is used to directly integrate the dimensionless forms of the above equations (see Appendix-A) in numerical simulations. The units of energy, length, and time are set by $\epsilon$, $\sigma$, $\tau_{u}=\sigma^{2}/\mu\epsilon$. The system property is controlled by the following dimensionless parameters: activity $\Lambda=v_{0}\sigma/\mu\epsilon$, translational diffusivity $\tilde{D}_{t}=D_{t}/\mu\epsilon$, and orientational diffusivity $\tilde{D}_{r}=D_{r}\sigma^{2}/\mu\epsilon$. In this paper, we focus on the impact of activity $\Lambda$, by setting $\tilde{D}_{r}=3$, $\tilde{D}_{t}=1$, and $\rho\sigma^{2}=1.1$, unless otherwise specified.
@@ III Non-affine parameter
The non-affine fluctuation $\chi_{0}$ at lattice position ${\mathbf{R}_{0}}$ is defined within a coarse-graining volume $\Omega$ containing neighbors $\{\mathbf{R}_{i}\}$ (see Fig. 1c). Instantaneous positions ${\mathbf{r}_{0}}$ and $\{\mathbf{r}_{i}\}$ yield displacements $\mathbf{u}_{0}=\mathbf{r}_{0}-\mathbf{R}_{0}$ and $\mathbf{u}_{i}=\mathbf{r}_{i}-\mathbf{R}_{i}$ relative to reference coordinates $\{\mathbf{R}_{i}\}$ on an ideal triangular lattice. The local strain matrix ${\cal E}$ is obtained by minimizing the mean-square difference $D^{2}$ between the observed relative displacements $\mathbf{\Delta}_{i}=\mathbf{u}_{i}-\mathbf{u}_{0}$ and estimated displacements ${\cal E}[\mathbf{R}_{i}-\mathbf{R}_{0}]$ [17],
$\chi_{0}=\sum_{i}\left(\mathbf{\Delta}_{i}-{\cal E}[\mathbf{R}_{i}-\mathbf{R}_{0}]\right)^{2}.$ (2)
Note that $D$ has the dimension of length. After minimization, the $\alpha\beta$-th component of the affine strain is given by
@@
where
$x_{\alpha\beta}$ $=\sum_{i=1}^{N_{\Omega}}\left(r_{i\alpha}-r_{0\alpha}\right)\left(R_{i\beta}-R_{0\beta}\right),$
$y_{\alpha\beta}$ $=\sum_{i=1}^{N_{\Omega}}\left(R_{i\alpha}-R_{0\alpha}\right)\left(R_{i\beta}-R_{0\beta}\right).$
The residual $D^{2}$ value using the minimized ${\cal E}$ denoted by ${\cal E}^{m}$ in Eq.(3) gives the non-affine fluctuation $\chi_{0}$ of the 00-th particle [17, 35]:
@@
A projection operator formalism [36] showed that this procedure projects the relative displacement field $\Delta_{i}$ onto mutually orthogonal subspaces of the affine strain tensor ${\cal E}^{m}$ and the non-affine parameter $\chi_{0}$. The global non-affine parameter $X=N^{-1}\sum_{i=1}^{N}\chi_{i}$ is the system average with $\chi_{i}$ denoting non-affine parameter associated with $i$-th particle. Note that $\chi$ and $X$ have dimensions of length squared, which we use in the next section to develop a scaling argument.
Figure 1: A schematic illustration of the activated solid. (a) An ideal triangular lattice that serves as the reference for calculation of non-affinity $\chi$. (b) A representative configuration at $\tilde{D}_{t}=1.0$ and $\Lambda=8.0$, is shown color coded by local non-affine parameter $\chi$; see the color bar. (c) A single coarse-graining region $\Omega$, which includes one central particle with reference position $\mathbf{R}_{0}$ (cyan) and actual location $\mathbf{r}_{0}$ (dark green). The reference (cyan, $\{\mathbf{R}_{i}\}$) and actual locations (dark green, $\{\mathbf{r}_{i}\}$) of its six nearest neighbors are shown as well.
@@ III.1 Scaling analysis
In an equilibrium harmonic solid with spring constant $k$ and temperature $k_{B}T$, the characteristic length scale of displacement fluctuations is $\ell^{2}=\frac{k_{B}T}{k}$, and the mean non-affine parameter is $\langle\chi\rangle_{eq}\sim\ell^{2}$ [36]. With translational diffusion $D_{t}=\mu k_{B}T$, similar scaling is expected in our system, where the elastic constant $k$, derivable from $U^{\prime\prime}(a)$, increases with density. The bulk, shear, and Young’s moduli of a two-dimensional solid increase non-linearly with density [77], predicting an associated decrease in $\langle\chi\rangle$. At melting, the non-affine parameter is expected to diverge as the shear modulus ${\cal G}$ vanishes. Thus, the expression can be generalized to $\langle\chi\rangle_{eq}\sim\frac{D_{t}}{{\cal G}}$, with ${\cal G}$ increasing with density $\rho$ and vanishing at the melting point. Ignoring non-linearities for simplicity, the following emerges:
@@
with $\rho_{m}$ representing the melting density.
In this paper, we first consider the impact of activity on a solid maintained at a fixed temperature and density. As activity increases, the persistence length $\ell_{p}=v_{0}/D_{r}$ dominates system’s behavior. The effective diffusivity of ABPs in two dimensions is $D_{\rm eff}=D_{t}+v_{0}^{2}/2D_{r}$ [78, 79], which includes the contribution from activity. The non-affine parameter is expected to increase with activity as $\langle\chi\rangle\sim D_{\rm eff}/{\cal G}$ which in dimensionless form is:
$\langle\chi\rangle={\cal G}^{-1}(\tilde{D}_{t}+\Lambda^{2}/2).$ (6)
In active systems, the melting point $\rho_{m}$ depends on $\Lambda$ [15], affecting the density dependence of ${\cal G}$ in Eq.(5). Beyond that, the shear modulus decreases with activity as
@@
as shown in a later section using numerical simulations. However, the $\Lambda$ dependence of ${\cal G}$ does not modify the dominant pre-melting scaling $\langle\chi\rangle\sim\Lambda^{2}$ appreciably.
In summary, for densities $\rho>\rho_{m}$ and activity $\Lambda<\Lambda_{c}$, the following scaling forms are expected:
$\langle\chi\rangle\sim(\rho-\rho_{m})^{-1}\quad\text{and}\quad\langle\chi\rangle\sim\Lambda^{2}.$ (8)

### v2

@@ Abstract
Internal activity can fundamentally reshape the mechanical behavior of solids, yet its role in softening and failure remains incompletely understood. In this study, we investigate spontaneous deformations in activated solids via non-affine fluctuations that quantify local rearrangements relative to global strain. Using scaling analysis and numerical simulations, we show that non-affinity in crystalline solids grows quadratically with active speed, increases linearly with persistence time before saturating, and scales inversely with the distance to the melting density. Spatial correlations reveal an activity-dependent growing correlation length, while relaxation dynamics are governed by the active persistence time. With increasing activity, the distributions of local non-affinity broaden, become more skewed, and develop heavy tails, eventually forming a secondary maximum that signals coexisting small and large non-affinities; this heterogeneity precedes defect formation and two-step melting from solid to hexatic and ultimately to fluid. Finally, we demonstrate that spatially patterned activation provides a simple route to locally induce non-affinity and mechanical softening. Our predictions are experimentally testable and suggest a pathway to tunable mechanics in adaptive metamaterials, with implications for mechanical regulation in biological systems.
@@ I Introduction
The study of active solids [1, 2, 3, 4, 5, 6] is inspired by living systems such as tissues and cell assemblies [7, 8, 9], bacterial biofilms [10, 11], and artificial material like smart matter [12, 13]. These systems are inherently out of equilibrium at the microscopic scale: their constituents locally consume and dissipate energy, generating stresses or propulsion and breaking detailed balance. Unlike crystalline solids, which break continuous translational and rotational symmetries in equilibrium, amorphous solids are stabilized by compression and by local force and torque balance [14, 15, 16, 17]. In active solids maintaining this balance is essential for mechanical stability; its breakdown naturally lead to deformation and flow.
Inspired by biological motility, artificial micro- and nano-swimmers have been developed that convert ambient energy into persistent motion and are well described by active Brownian particle dynamics [18, 19]. At high densities, such activity can profoundly alter the mechanical properties of solids. Recent studies have reported extreme deformations in active crystals [20] and yielding in active amorphous solids [21]. In equilibrium solids, low-energy deformations are captured by elastic displacement fields, while plasticity and yielding arise from defect-mediated, singular distortions [22, 23, 24, 25]. In amorphous solids, however, the absence of broken translational symmetry obscures this distinction, and localized, non-singular rearrangements play a significant role [26, 27, 28]. While homogeneous elasticity predicts a linear stress–strain relation, this description breaks down in the presence of disorder or spatially heterogeneous activity [29].
A natural framework to capture such effects is the decomposition of particle displacements into affine and nonaffine components [15, 30, 31]. The affine part describes uniform elastic distortions, whereas the nonaffine part encodes residual, disorder-driven rearrangements and defect precursors. In active solids, self-propulsion acts as a persistent, microscopic source of stress that enhances nonaffine motion even in the absence of external loading, making activity-induced nonaffinity a key ingredient in the mechanical response.
We investigate spontaneous deformations in active solids by quantifying local non-affine fluctuations relative to the global strain. Our main achievements are threefold. First, using scaling arguments and numerical simulations, we predict how the steady-state mean non-affinity depends on activity parameters – active speed and persistence – revealing behavior beyond simple active-temperature estimates. Second, at higher activity, pronounced non-affine zones coexist with regions of small non-affinity, preceding the divergence of non-affinity, defect proliferation, and eventual melting. Correlation functions show that the non-affine correlation length increases with active speed and grows to saturation with active persistence, while the relaxation of system-averaged non-affinity is set by the active persistence time. Third, we demonstrate that spatially patterned activation provides a direct route to locally induce non-affinity and mechanical softening. Together, these results establish a quantitative framework for understanding and controlling the interplay between activity, non-affine deformations, and mechanical stability in dense active materials.
The paper is organized as follows. In Sec. II, we introduce the model of the active solid. Sec. III presents the calculation of non-affine fluctuations, the steady-state scaling relations, and the local distributions and correlations of non-affinity. Sec. IV examines changes in structural indicators, including the shear modulus, solid and hexatic order parameters, and topological defect proliferation. Sec. V demonstrates how spatially patterned activity can locally induce non-affinity and soften the solid. Finally, Sec. VI summarizes the main results and provides a discussion and outlook.
@@ II Model
Modeling the constituents of this high-density solid ($\rho\sigma^{2}\geq 1$) as Active Brownian Particles (ABPs), we investigate its properties under increasing activity, controlled by the active speed and persistence. Each particle moves with self-propulsion speed $v_{0}$ in a direction $\hat{\mathbf{n}}_{i}={\cos\theta_{i}(t),\sin\theta_{i}(t)}$, with the angle $\theta_{i}(t)$ evolving via rotational diffusion. The dynamics of the $i$-th particle is governed by the following coupled Langevin equations: {align} d r_i(t) = v_0^n_i(t) dt - μ∇_i∑_j ϵR_i U(r_ij) dt + 2D_t  dB_i(t)
dθ_i(t) = 2D_r dB^r_i(t), where $D_{t}$ and $D_{r}$ are the translational and rotational diffusivities, and $d\mathbf{B}_{i}(t)$ and $dB_{i}^{r}(t)$ are independent Gaussian white noise terms with mean zero and variance $dt$. The Euler-Maruyama method is used to directly integrate the dimensionless forms of the above equations (see Appendix-A) in numerical simulations. The units of energy, length, and time are set by $\epsilon$, $\sigma$, $\tau_{u}=\sigma^{2}/\mu\epsilon$. The system property is controlled by the following dimensionless parameters: active speed $\Lambda=v_{0}\sigma/\mu\epsilon$, translational diffusivity $\tilde{D}_{t}=D_{t}/\mu\epsilon$, and orientational diffusivity $\tilde{D}_{r}=D_{r}\sigma^{2}/\mu\epsilon$. Unless stated otherwise, we set $\tilde{D}_{t}=1$ and $\tilde{D}_{t}=0$, to simulate thermal or athermal systems, respectively, and fix the density at $\rho\sigma^{2}=1.1$. We also examine the effects of activity by varying $\Lambda$} and $\tilde{D}_{r}$, as well as the effect of changing density $\rho\sigma^{2}$.
@@ III Non-affine parameter
To characterize deviations from uniform, affine deformations in active solids, we define a non-affinity parameter, which measures the extent of local rearrangements relative to a global strain. Specifically, for a given local region, it quantifies the discrepancy between the observed particle displacements and those predicted by a best-fit affine transformation.
The non-affine fluctuation $\chi_{0}$ of the 0-th particle, at lattice position $\mathbf{R}_{0}$}, is defined within a coarse-graining volume $\Omega$ containing all six nearest neighbors $\{\mathbf{R}_{i}\}$ of $\mathbf{R}_{0}$} (see Fig. 1c). Instantaneous positions $\mathbf{r}_{0}$} and $\{\mathbf{r}_{i}\}$ yield displacements $\mathbf{u}_{0}=\mathbf{r}_{0}-\mathbf{R}_{0}$ and $\mathbf{u}_{i}=\mathbf{r}_{i}-\mathbf{R}_{i}$ relative to reference coordinates $\{\mathbf{R}_{i}\}$ on an ideal triangular lattice.
[…]
$D^{2}=\sum_{i}\left(\mathbf{\Delta}_{i}-{\cal E}[\mathbf{R}_{i}-\mathbf{R}_{0}]\right)^{2}.$ (1)
Note that $D$ has the dimension of length. After minimizing $D^{2}$ at a given density and activity, one obtains the $\alpha\beta$-th component of the optimal strain matrix, the affine strain,
@@
where {align} x_αβ = ∑_i=1^N_Ω ( r_i α - r_0 α ) ( R_i β - R_0 β ),
y_αβ = ∑_i=1^N_Ω ( R_i α - R_0 α ) ( R_i β - R_0 β ).
[…]
@@
Applying this method to all particles results in the local non-affinity $\chi_{i}$ for each particle $i$ in the system. A projection operator formalism showed that this procedure projects the relative displacements $\mathbf{\Delta}_{i}$ onto mutually orthogonal subspaces of the affine and non-affine fluctuations, characterized by the affine strain tensor ${\cal E}^{m}$ and the scalar non-affine parameter $\chi_{i}$ [31].
The global non-affine parameter $X=N^{-1}\sum_{i=1}^{N}\chi_{i}$ is the system average with $\chi_{i}$ denoting non-affine parameter associated with $i$-th particle. Note that $\chi_{i}$ and $X$ have dimensions of length squared, which we use in the next section to develop a scaling argument.
Figure 1: A schematic illustration of the activated solid. (a) An ideal triangular lattice that serves as the reference for the calculation of non-affinity $\chi$. (b) A representative configuration at $\tilde{D}_{t}=1.0$ and $\Lambda=8.0$, is shown color coded by local non-affine parameter $\chi$; see the color bar. (c) A single coarse-graining region $\Omega$ is defined as the volume enclosed by the first nearest neighbors, determined by mean inter-particle separation $a$ at a given denisty $\rho$. This region includes one central particle with reference position $\mathbf{R}_{0}$ (cyan) and actual location $\mathbf{r}_{0}$ (dark green). The reference (cyan, $\{\mathbf{R}_{i}\}$) and actual locations (dark green, $\{\mathbf{r}_{i}\}$) of its six nearest neighbors are shown as well.
@@ III.1 Scaling analysis
In an equilibrium solid with harmonic relaxation rate $\mu k$ and single particle diffusivity $D_{t}$, the characteristic length scale associated with particle rearrangements, $\ell_{\rm eq}$, is set by a balance of diffusion and relaxation, yielding $\ell_{\rm eq}^{2}\sim D_{t}/\mu k$. For an equilibrium harmonic solid, the $n$-th moment of the mean non-affine parameter, $X=N^{-1}\sum_{i}^{N}\chi_{i}$, was shown to scale as $\langle X^{n}\rangle_{eq}\sim\ell^{2n}$ [31]. In sterically stabilized solids, the effective elastic constant $k$ is determined by $U^{\prime\prime}(a)$, and therefore increases with density. For a two-dimensional isotropic solid, the shear and Young’s moduli scale linearly with $k$ and, more generally, increase nonlinearly with density [35]. This elastic stiffening implies a corresponding decrease in the mean non-affinity, $\langle X\rangle$. Approaching melting, however, the shear modulus $\cal G$} vanishes, and the non-affine parameter is expected to diverge. These considerations motivate a generalized expression for equilibrium non-affine rearrangements, $\langle X\rangle_{eq}\sim\frac{D_{t}}{{\cal G}}$, with $\cal G$} increasing with density $\rho$ and vanishing at the melting transition. Ignoring non-linearities for simplicity, the density dependence of shear modulus can be approximated as:
@@
In this paper, we first consider the impact of activity on a dense solid. As activity increases, the persistence length $\ell_{p}=v_{0}/D_{r}$ begins to play an important role. One might then expect that replacing the thermal diffusivity $D_{t}$ by the effective active diffusivity $v_{0}^{2}/2D_{r}$ [36, 37] in the expression $D_{t}/{\cal G}$ would be sufficient to account for the excess non-affinity induced by activity. However, this simple substitution fails in a solid, where each test particle is trapped within a cage formed by its neighbors. Approximating this confinement by a harmonic potential with stiffness $k$, we can invoke an earlier estimate for the steady-state mean-squared displacement arising from activity $\ell_{\rm ac}^{2}=v_{0}^{2}/[\mu k(D_{r}+\mu k)]$, as obtained in Ref. [38]. This leads us to hypothesize that activity induces an excess non-affinity of the form $\langle X^{(a)}\rangle\approx v_{0}^{2}/[\alpha{\cal G}(D_{r}+\alpha{\cal G})]$, where the spring constant is replaced by the more appropriate solid descriptor, the shear modulus $\cal G$}; here $\alpha$ is a proportionality constant. Thus, the total non-affine parameter can be written in the following dimensionless form as
$\frac{\langle X\rangle}{\sigma^{2}}\approx\frac{\tilde{D}_{t}}{\alpha\tilde{\cal G}}+\frac{\Lambda^{2}}{\alpha{\tilde{\cal G}}(\tilde{D}_{r}+\alpha\tilde{\cal G})}$. (5)
In the above expression, we used dimensionless forms of translational diffusivity $\tilde{D}_{t}=D_{t}/\mu\epsilon$, shear modulus $\tilde{\cal G}={\cal G}\sigma^{2}/\epsilon$, rotational diffusivity $\tilde{D}_{r}=D_{r}\sigma^{2}/\mu\epsilon$ and active speed $\Lambda=v_{0}\sigma/\mu\epsilon$.
[…]
@@
as demonstrated in a later section through numerical simulations (Fig. 6). However, a small value of ${\cal G}_{1}$ ensures that the $\Lambda$ dependence of $\cal G$} does not modify the dominant pre-melting scaling of the excess non-affinity due to activity $\langle X^{(a)}\rangle=\langle X\rangle-\langle X\rangle_{eq}\sim\Lambda^{2}$ appreciably.
In summary, for densities $\rho>\rho_{m}$ and active speed $\Lambda<\Lambda_{c}$ allowing the system to remain in the solid phase, the following scaling forms for $\langle X^{(a)}\rangle$ are expected:
$\langle X^{(a)}\rangle\sim(\rho-\rho_{m})^{-1}\quad\text{and}\quad\langle X^{(a)}\rangle\sim\frac{\Lambda^{2}}{\tilde{\cal G}[\tilde{D}_{r}+\alpha\tilde{\cal G}]}$. (7)